Doede Corporation uses activity-based costing to compute product margins. In the first stage, the activity-based costing system allocates two overhead accounts--equipment depreciation and supervisory expense--to three activity cost pools--Machining, Order Filling, and Other--based on resource consumption. Data to perform these allocations appear below:

Overhead costs:

Equipment depreciation $ 35,000
Supervisory expense $ 12,000

Distribution of Resource Consumption Across Activity Cost Pools:

Activity Cost Pools
Machining Order Filling Other
Equipment depreciation 0.40 0.30 0.30
Supervisory expense 0.40 0.20 0.40

In the second stage, Machining costs are assigned to products using machine-hours (MHs) and Order Filling costs are assigned to products using the number of orders. The costs in the Other activity cost pool are not assigned to products.

Activity:

MHs
(Machining)
Orders
(Order Filling)
Product W1 5,660 109
Product M0 22,100 932
Total 27,760 1,041

Finally, sales and direct cost data are combined with Machining and Order Filling costs to determine product margins.

Sales and Direct Cost Data:

Product W1 Product M0
Sales (total) $ 64,150 $ 60,700
Direct materials (total) $ 33,400 $ 22,600
Direct labor (total) $ 17,200 $ 32,600

What is the overhead cost assigned to Product W1 under activity-based costing? (Round your intermediate calculations to 2 decimal places and your final answer to nearest whole dollar amount.)
Answer #1

Answer

  • Workings

Machining

Order Filling

Other

Total

Equipment depreciation

$14,000

$10,500

$10,500

$35,000

Supervisory expense

$4,800

$2,400

$4,800

$12,000

A

Total

$18,800

$12,900

$15,300

B

Total Allocation base

27760

1041

C = A/B

Allocation rate

$0.68

$12.39

--Calculations

Machining

Order Filling

Other

Equipment depreciation

=35000*0.4

=35000*0.3

=35000*0.3

Supervisory expense

=12000*0.4

=12000*0.2

=12000*0.4

  • Answer

Overhead cost assigned to Product W1:

Machining cost

$3,849

[5660 MH x $ 0.68]

Order Filling Cost

$1,351

[109 orders x $ 12.39]

Total

$5,199

Answer